# Recognition of chordal graphs and cographs which are Cover-Incomparability graphs

### v3

@@ Abstract
Key Words: Cover-Incomparability graphs; Chordal graph; Cographs; Recognition Algorithm.
@@ 3.1 Recognition of chordal C-I graphs
  • Step 1:
    Apply the perfect elimination ordering (PEO) on $G$ and check whether $G$ is a chordal graph or not. If there is no PEO then return $G$ not chordal and stop. Otherwise, go to Step 2.
  • Step 2:
    Using the PEO find the clique tree $T$ of the given graph $G$ using Blair-Peyton algorithm. If the clique tree has more than 2 leaf nodes, then return $G$, not chordal C-I graph, and stop. Otherwise, go to Step 3. (Now the clique tree is a path, and the vertices are in order $C_{1},C_{2},\ldots,C_{k}$)
  • Step 3:
    Check whether there is an element in $C_{i}$ that is not in $C_{i+1}$ and $C_{i-1}$ for $i=2,3,\ldots,k-1$. If such an element exists at any stage, then stop and return $G$ as not a chordal C-I graph. Otherwise, return $G$ as a chordal C-I graph.
@@ 4 Structure of C-I cographs
  • Step 2:
    If BFS completes in one step by reaching all the leaf nodes ( if all the adjacent nodes of the root node are leaf nodes), then the graph is a complete graph and it is C-I cograph; stop, and we are done. Otherwise, go to Step 3.
@@
    After Step 2, only 0-nodes are obtained. Then $G$ is not a C-I cograph and Stop.
  • Step 3(ii):
    After Step 2, both 0-nodes and leaf nodes are obtained. Check whether the number of zero nodes is greater than the number of leaf nodes. If yes, then $G$ is not a C-I cograph and stop. Otherwise goto Step 4
@@
    Continue the searching at level 2. If 00-nodes has more than two children as $1$-nodes then $G$ is not a C-I cograph and stop. Otherwise, go to Step 5.
  • Step 5:
    Continue searching at level 3. In level 3, if any $1$-node has 00-node as a child, then $G$ is not a C-I cograph and stop. If all the children of the $1$-nodes are the leaf node then we are done. the resulting graph is a C-I cograph.
  • Step 6:
    Label the leaf nodes of each 0-node  branch in level 2 as $C_{1}^{i}$ and $C_{3}^{i}$. Partition the leaf nodes from the rooted node 1 into the number of 0-nodes. That is, the leaf nodes are partitioned into $r$ sets, $C_{2}^{i}$ for $i=1,2,\ldots,r$, where $r$ denotes the number of 0-nodes at level 2.
  • Step 7:
    Construct completely ranked posets $P_{i}$ with rank 0 elements as $C_{1}^{i}$, rank 1 elements as $C_{2}^{i}$, and rank 2 elements as $C_{3}^{i}$ for $i=1,2,3,\ldots,r$. Return $P=P_{1}+P_{2}+\cdots+P_{r}$ as one of the poset of the C-I cograph $G$.
@@ Acknowledgment:
Anil, A. acknowledges the financial support from the University of Kerala, for providing University Post Doctoral Fellowship (Ac EVII 5576/2023/UOK dated 30/06/2023).

### v4

@@ Abstract
keywords:
Cover-Incomparability graphs; Chordal graph; Cographs; Recognition Algorithm.
@@ 3.1 Recognition of chordal C-I graphs
  • 1Step 1:
    Apply the perfect elimination ordering (PEO) on $G$ and check whether $G$ is a chordal graph or not.
If there is no PEO then return $G$ not chordal and stop. Otherwise, go to Step 2.
  • 2Step 2:
    Using the PEO find the clique tree $T$ of the given graph $G$ using Blair-Peyton algorithm.
  • 3 If the clique tree has more than 2 leaf nodes, then return $G$, not chordal C-I graph, and stop.
    Otherwise, go to Step 3. (Now the clique tree is a path, and the vertices are in order $C_{1},C_{2},\ldots,C_{k}$)
  • 4Step 3:
    Check whether there is an element in $C_{i}$ that is not in $C_{i+1}$ and $C_{i-1}$ for $i=2,3,\ldots,k-1$.
  • 5 If such an element exists at any stage, then stop and return $G$ as not a chordal C-I graph.
    Otherwise, return $G$ as a chordal C-I graph.
@@ 4 Structure of C-I cographs
    • 1Step 2:
      If BFS completes in one step by reaching all the leaf nodes (if all the adjacent nodes of the root node are
    2 leaf nodes), then the graph is a complete graph and it is C-I cograph; stop, and we are done.
    Otherwise, go to Step 3.
@@
    After Step 2, only 0-nodes are obtained. Then $G$ is not a C-I cograph and Stop. 3Step 3(ii): After Step 2, both 0-nodes and leaf nodes are obtained. Check whether the number of zero nodes is
  • 4 greater than the number of leaf nodes. If yes, then $G$ is not a C-I cograph and stop.
    Otherwise goto Step 4
@@
    Continue the searching at level 2. If 00-nodes has more than two children as $1$-nodes then $G$ is not a C-I cograph and stop. Otherwise, go to Step 5. 5Step 5: Continue searching at level 3. In level 3, if any $1$-node has 00-node as a child, then $G$ is not a C-I cograph
  • and stop. If all the children of the $1$-nodes are the leaf node then we are done. The resulting graph is a C-I cograph.
  • 6Step 6:
    Label the leaf nodes of each 0-node  branch in level 2 as $C_{1}^{i}$ and $C_{3}^{i}$. Partition the leaf nodes from
  • the rooted node 1 into the number of 0-nodes. That is, the leaf nodes are partitioned into $r$ sets, $C_{2}^{i}$ for $i=1,2,\ldots,r$, where $r$ denotes the number of 0-nodes at level 2.
  • 7Step 7:
    Construct completely ranked posets $P_{i}$ with rank 0 elements as $C_{1}^{i}$, rank 1 elements as $C_{2}^{i}$, and rank 2 elements as $C_{3}^{i}$ for $i=1,2,3,\ldots,r$.
  • Return $P=P_{1}+P_{2}+\cdots+P_{r}$ as one of the poset of the C-I cograph $G$.
@@ Acknowledgment:
    Arun Anil acknowledges the financial support from the University of Kerala, for providing University Post Doctoral Fellowship (Ac EVII 5576/2023/UOK dated 30/06/2023).